OpenStudy (anonymous):

Find the volume of a cube with the given side length. 10 units A. 10,000 cubic units B. 1000 cubic units C. 100 cubic units D. 30 cubic units

OpenStudy (kidrah69):

\(\huge\color{orange}{WELCOME~TO~OPENSTUDY!!~:)}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://goo.gl/5pp1u0 For an image tutorial look at this! Its amazing :o https://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/ Also consider reading the CoC and the T&C http://openstudy.com/code-of-conduct http://openstudy.com/terms-and-conditions

OpenStudy (kidrah69):

Since we are talking about a cube we want to cube the side length heres the formula \(\huge\color{red}{V=a^3}\)

OpenStudy (kidrah69):

Our side lenth is to so we plug it to where the A is \(\huge\color{green}{V=(10)^3}\)

OpenStudy (anonymous):

The answer would be B. Because a cube uses 4 edges, so you would do, 10 * 10 * 10 * 10=?